# Optical and UV Flares from Binary Black Hole Mergers in Active Galactic Nuclei

### v1

@@ Appendix A The initial pressure of the plasma ejection
We estimate a lower limit for the typical values of the parameter $A$ using $P\approx E_{\mathrm{0,th}}/V_{0}$, which we underestimate using $V_{0}\approx(3\pi/8)h_{\mathrm{d}}^{3}/\theta_{\mathrm{k}}^{2}$, where $V_{0}$ is an overestimation for the initial volume of the disc eruption (see equation 21), and $E_{\mathrm{0,th}}=E_{\mathrm{0,ll}}/2$ given by equation 12 is an underestimation for $E_{0}$. Then we obtain:
$A_{\mathrm{ll}}$ $\sim\frac{16}{81}\frac{a_{\mathrm{r}}m_{\mathrm{p}}^{3}}{k_{\mathrm{B}}^{4}}\frac{(\eta_{\mathrm{w}}\eta_{\mathrm{in}})^{3}}{n_{\mathrm{i}}}\left(\frac{\theta_{\mathrm{k}}GM_{\bullet}c}{h_{\mathrm{d}}}\right)^{6}\frac{1}{v_{\mathrm{k}}^{12}}$ (50)
$\sim 4.3\times 10^{6}\left(\frac{\eta_{\mathrm{w}}}{0.005}\right)^{3}\left(\frac{\eta_{\mathrm{in}}}{0.1}\right)^{3}\left(\frac{10^{\mathrm{14}}\mathrm{cm}^{-3}}{n_{\mathrm{i}}}\right)$
×$\times\left(\frac{\theta_{\mathrm{k}}}{0.1}\right)^{6}\left(\frac{M_{\bullet}}{40\mathrm{M}_{\odot}}\right)^{6}\left(\frac{0.01}{h_{\mathrm{d}}/a}\right)^{6}\left(\frac{9000}{a/R_{\mathrm{g}}}\right)^{6}$
×$\times\left(\frac{10^{6}\mathrm{M}_{\odot}}{M_{\mathrm{S}}}\right)^{6}\left(\frac{400\mathrm{kms}^{-1}}{v_{\mathrm{k}}}\right)^{12},$

### v2

@@ Appendix B The initial pressure of the plasma ejection
We estimate a lower limit for the typical values of the parameter $A$ using $P\approx E_{\mathrm{0,th}}/(3V_{0})$, which we underestimate using $V_{0}\approx\pi h_{\mathrm{d}}^{3}/(2\theta_{\mathrm{k}}^{2})$, where $V_{0}$ is an over-estimation for the initial volume of the disc eruption (see equation 21), and $E_{\mathrm{0,th}}=E_{\mathrm{0}}/2$, where $E_{0}$ is given by equation 15. Then we obtain:
$A$ $\sim\frac{1}{162}\frac{a_{\mathrm{r}}m_{\mathrm{p}}^{3}}{k_{\mathrm{B}}^{4}}\frac{(\eta_{\mathrm{w}}\eta_{\mathrm{in}})^{3}}{n_{\mathrm{i}}}\left(\frac{\theta_{\mathrm{k}}GM_{\bullet}c}{h_{\mathrm{d}}}\right)^{6}\frac{1}{v_{\mathrm{k}}^{12}}$ (64)
$\sim 8.229\times 10^{6}\left(\frac{\eta_{\mathrm{w}}}{0.05}\right)^{3}\left(\frac{\eta_{\mathrm{in}}}{0.1}\right)^{3}\left(\frac{10^{\mathrm{14}}\,\mathrm{cm}^{-3}}{n_{\mathrm{i}}}\right)$
×$\times\left(\frac{\theta_{\mathrm{k}}}{10^{\circ}}\right)^{6}\left(\frac{M_{\bullet}}{20\mathrm{M}_{\odot}}\right)^{6}\left(\frac{0.005}{h_{\mathrm{d}}/a}\right)^{6}\left(\frac{10^{4}}{a/R_{\mathrm{g}}}\right)^{6}$
×$\times\left(\frac{10^{7}\,\mathrm{M}_{\odot}}{M_{\mathrm{S}}}\right)^{6}\left(\frac{200\,\mathrm{kms}^{-1}}{v_{\mathrm{k}}}\right)^{12},$